# Quantum Adaptive Self-Attention for Quantum Transformer Models

### v2

@@ 7 Conclusion
In this work, we introduced Quantum Adaptive Self-Attention (QASA), a novel hybrid attention mechanism that integrates parameterized quantum circuits into the Transformer architecture to enhance temporal sequence modeling. By replacing classical dot-product attention with a QNN-based adaptive mechanism and incorporating a residual quantum feature refinement layer, QASA demonstrates significant improvements in convergence speed, generalization, and predictive accuracy, particularly in resource-constrained scenarios such as time-series forecasting. Empirical evaluations on synthetic benchmarks confirm the model’s ability to capture complex temporal dependencies with fewer layers and lower validation loss compared to classical baselines. Furthermore, we provide preliminary complexity analysis suggesting a potential quantum advantage in gradient computation, supported by observations aligned with theoretical expectations under quantum complexity assumptions. These results underscore the promise of quantum-enhanced architectures in next-generation deep learning and open new directions for future research in scalable, hardware-compatible quantum neural models.
Acknowledgements EJK thanks National Yang Ming Chiao Tung University for its support.

### v3

@@ 5 Conclusion
In this work, we introduced Quantum Adaptive Self-Attention (QASA) and, through systematic experimentation, established an architectural parsimony principle for hybrid quantum-classical design. Our findings converge on a single insight: the optimal strategy for quantum integration is not to maximize quantum resources, but to place minimal quantum computation where it matters most.
This principle rests on four empirical pillars:
  1. 1.
    One layer is enough. Our ablation study demonstrates that a single quantum layer at the optimal position outperforms all multi-layer configurations. Increasing from 1 to 4 quantum layers degrades performance on most tasks, and the optimal position is task-dependent (last layer for chaotic dynamics, penultimate for smooth oscillations).
  2. 2.
    Minimal resources, maximal impact. With only 36 quantum parameters and 27 CNOT gates, QASA achieves near-maximal entangling capability ($Q=0.981$) and matches or exceeds QLSTM (128 params, 56 CNOTs) and QnnFormer (90 params) across nine benchmarks. QASA achieves the best MSE on four tasks, significantly outperforming QLSTM on seasonal trend ($p=0.009$, Cohen’s $d>6$).
  3. 3.
    Noise-resilient, not noise-brittle. Under depolarizing noise at levels typical of current hardware ($p=0.01$), QASA’s performance remains stable or even improves (MAE $10.8\%$ on chaotic logistic, confirmed across 3 seeds). While the noise-induced regularization effect observed on a single task requires further validation across additional tasks and noise channels, it demonstrates that QASA does not require error-corrected hardware for deployment on NISQ devices.
  4. 4.
    Fewer layers avoid barren plateaus. Gradient variance analysis reveals that increasing QASA from 1 to 2+ entangling layers causes a $30\times$ drop in gradient magnitude, approaching a barren plateau. The single-layer design is not merely sufficient—it is necessary to maintain trainability.
Beyond these architectural insights, our work contributes a task-conditional taxonomy: quantum enhancement benefits chaotic, noisy, and trend-dominated signals (chaotic logistic, seasonal trend, ETTh1 with 6.0% MAE reduction), while classical Transformers remain superior for clean periodic waveforms (damped oscillator, d=1.18d=-1.18; sawtooth, d=0.99d=-0.99). This taxonomy provides practitioners with a principled guide for when to deploy—and when not to deploy—quantum-enhanced models.
Our preliminary complexity analysis further suggests a potential reduction in gradient computation lower bounds from $\Omega(T^{2})$ to $\Omega(T)$ for quantum attention mechanisms, though a rigorous proof remains an open problem.
Future work should focus on: (1) adaptive hybrid strategies that selectively engage quantum layers based on detected signal characteristics, (2) validation on additional real-world time-series datasets at larger scale, (3) systematic study of noise-induced regularization across different noise channels and circuit depths, and (4) formal theoretical analysis of the conjectured quantum speedup in attention gradient computation.
[…]
EJK acknowledges financial support from the National Science and Technology Council (NSTC) of Taiwan under Grant No. NSTC 114-2112-M-A49-036-MY3.
Author contributions
C.-S.C.: conceptualization, methodology, software, writing—original draft. E.-J.K.: supervision, theoretical analysis, writing—review & editing.
Competing interests
The authors declare no competing interests.
Data availability statement
The data and source code that support the findings of this study are openly available at https://github.com/ChiShengChen/QASA. The ETTh1 benchmark used for real-world validation is publicly available at https://github.com/zhouhaoyi/ETDataset. All synthetic datasets can be regenerated from the equations in Section 4.2 using the fixed seeds listed in Section 4.10.